# Energy Weighted Learning Progress Guided Interleaved Multi-Task Learning

### v1

@@ IV-B Task Arbitration
0:  Number of tasks $T$, number of epochs $K$, exploration rate $\epsilon=0.1$
  for epoch $=1$ to $K$ do
     Compute score $s^{(t)}$ for each task $t\in\{1,2,\ldots,T\}$
@@
     Generate a random number $r$ uniformly from $[0,1]$
@@
        $t^{*}\leftarrow$ Random task from $\{1,2,\ldots,T\}\setminus\{t_{\text{winner}}\}$
@@
        $t^{*}\leftarrow t_{\text{winner}}$
@@
     Perform training on the selected task $t^{*}$
@@
    Learning Progress (LP) Based Task Selection Similar to LP definition in[11], in our work LP is assessed by examining the recent evolution of a task’s error signal. For a given task $t$, let $E^{(t)}_{i}$ be the error at time step $i$. We track the error over the last $L$ training steps (e.g., $L=5$) to compute the slope of the error curve. Consider the set of errors $\{E^{(t)}_{i-L+1},E^{(t)}_{i-L+2},\ldots,E^{(t)}_{i}\}$. We fit a linear model to these points to estimate the slope $\beta^{(t)}$. If $\beta^{(t)}<0$, the error is decreasing, and the task is making positive learning progress. If $\beta^{(t)}\geq 0$, it suggests that the error has plateaued or is increasing, implying no recent improvement. Under the LP-based interleaving scheme, we compute the learning progress $LP^{(t)}$ for each task $t$. We then select the task with the highest LP (most improving) to be trained next. In other words, tasks that are currently showing rapid improvement receive more training time, while tasks that have stalled or regressed receive less immediate attention. Formally, if $\beta^{(t)}$ is the slope for task $t$, then we choose $t^{*}$ in two steps:
@@
      ${LP^{(t)}}=\begin{cases}\lvert\beta^{(t)}\rvert,&{\text{if }}\ \beta^{(t)}<0\\ {0,}&{\text{otherwise.}}\end{cases}$
@@
    Energy Consumption (EC) Modulated Learning Progress (LP) Based Task Selection. While learning progress alone can guide the model toward tasks showing improvement, it does not consider the computational or energetic cost associated with training each task. In scenarios where energy efficiency is a concern, we incorporate energy consumption into the interleaving strategy. We define the neurons’ activation “energy” consumption $EC^{(t)}$ for task $t$ as the cumulative energy used over the last $L$ training steps. This could be measured in terms of computational operations, memory usage, or any proxy for energy expenditure. Lower $EC^{(t)}$ values mean the task has recently been trained efficiently, while higher values indicate that the task has been relatively costly in terms of energy. However, using EC only as the interleaving guidance, the network will only choose the tasks that produce low activation energy without taking the learning performance of the tasks into consideration, which can lead to worsened overall multi-task learning performance (see Figure 7a). To blend both learning performance and energy considerations, we employ the following combined score for each task $t$:
    $s^{(t)}=\exp{(k\cdot LP^{(t)})}/EC^{(t)},$ (9)
    where:
    • $LP^{(t)}$ is the learning progress defined above.
    • $EC^{(t)}$ is recent energy consumption for the task.
    • $k$ is a positive constant that controls the sensitivity of the combined score to energy consumption ($EC^{(t)}$).
    Note that, to combine LP and EC in a balanced manner, we first normalize both values between $[0,1]$ before computing the combined score. The $s^{(t)}$ can be interpreted as follows:
    • When $k$ approaches to zero, $\exp{(k\cdot LP^{(t)})}\approx 1$, thus the score is primarily influenced by $EC^{(t)}$, even if the task has strong LP. This discourages spending too many resources on a single, costly task.
    • When $k$ is high, $\exp{(k\cdot LP^{(t)})}$ becomes dominant, and thus EC becomes negligible.
    Under the EC-modulated LP-based interleaving strategy, we compute combined scores for each task $t$ as given above. The next task to train is the one with the highest combined score:
@@
    This ensures a balance between quick gains in task performance and maintaining overall energy efficiency. Tasks with high learning progress but also low recent energy costs are favored, while tasks that have become too energy-intensive or are not showing improvement receive less immediate focus. By integrating these interleaving strategies, we encourage a more dynamic, human-like learning schedule. The LP-based method shifts focus toward tasks that are currently “improvable,” while the EC-modulated LP method adds a layer of energy awareness, guiding the interleaving process toward both efficiency and sustained improvement. An overview of the task arbitration mechanism used in this work is presented in Algorithm 1, where a non-greedy selection strategy is adopted by giving other tasks a chance to be selected, with an exploration rate $\epsilon$, thereby reducing over-reliance on the highest-scoring task.
IV-C Simulation Environment
@@
In our study, we use a simulated tabletop environment to evaluate multi-task learning through effect prediction tasks. The simulation is implemented using PyBullet engine [90] and utilizes a UR10 robotic arm equipped with a Franka Panda end effector. The tabletop is a constrained environment designed to replicate real-world conditions for object interaction tasks. A variety of objects with differing shapes, sizes, and physical properties are placed on the table for the robot to interact with during the tasks. The objects used in this study include spheres, cubes, cylinders, and square prisms. Cylinder and square prism can be placed either horizontally or vertically, resulting in a total of six distinct object types.
IV-D Effect Prediction Tasks
In this work, we define three action-effect prediction tasks performed in the simulated environment explained above. Each task requires the model to predict the resulting state of one or more objects after a specific action is executed. The effect prediction tasks are designed as follows:
@@ IV-D Effect Prediction Tasks
Push task. A single object is randomly placed on the table. The robot applies a “pushing” action with its end effector to the object’s center of mass (CoM), with an angle chosen between $[0,180]$ degrees. The state of the object includes its Cartesian position (x, y, z) and orientation, represented using the sine and cosine of each Euler angle (x, y, z axes), resulting in a 9-dimensional state representation:
$\mathbf{s}=[x,y,z,\sin(\phi_{x}),\cos(\phi_{x}),$
$\sin(\phi_{y}),\cos(\phi_{y}),\sin(\phi_{z}),\cos(\phi_{z})].$
@@
Hit Task. The second task is a dynamic variation of the Push task. The setup is the same, but the pushing action is applied with twice the velocity, resulting in more chaotic and less predictable object behavior. The state representation and action encoding remain identical to the Push task.
@@
Stack Task. In the final task, the goal is to place one object on top of another, forming a possible stable stack. This task involves two objects: a moving object and a target object. As a result, the state includes the positional and orientation information for both objects, doubling the state information compared to the Push and Hit tasks.
$\mathbf{s}=[\mathbf{s}_{m},\mathbf{s}_{t}],$
@@
where $\mathbf{s}_{m}$ and $\mathbf{s}_{t}$ are the 9-dimensional states of the moving and target objects, respectively. The action is represented by concatenating the one-hot encodings of both the moving object and the target object:
$\mathbf{a}=[\text{onehot}(o_{m}),\text{onehot}(o_{t})].$
V Experiments and Results
To evaluate the effectiveness of our proposed method, which utilizes EWLP-based task selection within an interleaved multi-task learning framework, we perform a series of experiments across different aspects. The central aim of these experiments is to investigate how intelligent task interleaving influences learning efficiency, knowledge transfer, and final task performance in comparison to other approaches. We begin by comparing INTER-LP to several baselines designed to isolate the contributions of individual components. These include a single-task learning setup (SINGLE), a multi-task model with random task selection (INTER-RAND), and a blocked learning setup where tasks are trained sequentially in fixed permutations (BLOCK). In addition to these, we conduct a set of experiments on our energy-aware variant, INTER-LPE, which extends INTER-LP by incorporating energy consumption into the task arbitration mechanism. These experiments are designed to explore how varying the energy sensitivity coefficient $k$ affects the trade-off between prediction performance and computational cost. By analyzing different values of $k$, we demonstrate the flexibility of INTER-LPE in adapting to resource constraints while maintaining competitive learning outcomes. Since we maintain a consistent model architecture across all settings, we ensure that any observed performance differences are attributable to task arbitration strategies rather than architectural complexity or capacity. The experiments are structured to evaluate overall learning performance across training regimes, investigating task-wise learning dynamics to examine how individual tasks benefit from shared training and interleaving, assessing robustness to network complexity by comparing models across low, medium, and high parameter counts, demonstrating the advantages of interleaved learning over blocked training in mitigating catastrophic forgetting, and conducting ablation studies to reveal the individual and joint contributions of architectural components such as the shared attention mechanism and task-specific flag bits. Together, these evaluations provide a comprehensive picture of how and why EWLP-guided interleaved learning can outperform traditional scheduling strategies in multi-task environments.
V-A Baselines
To demonstrate the advantages of our interleaved multitask learning method with LP-based task selection (INTER-LP), we compare it against the following baselines:
@@ V-A Baselines
  • INTER-RAND: The second baseline uses the shared architecture used in our proposed method, without any modification but removes the LP-based task selection mechanism. Instead, tasks are chosen randomly at each training iteration, with each task having an equal chance of being selected. This baseline serves to highlight the impact of our task selection (interleaving) strategy by comparing it against a scenario where the interleaving is uniformly random. This baseline helps isolate the effectiveness of the scheduling algorithm from the benefits of shared parametrization.
@@
  • BLOCK: The third and final baseline is implemented using the same shared network architecture but trains tasks in dedicated, uninterrupted blocks rather than interleaving them. To account for any potential effects caused by the order in which tasks are learned, we evaluate all six possible permutations of the three tasks (denoted as BLOCK-xyz, where the xyz represents the order of task training with their corresponding ids). In our case, we have three tasks, namely Push (0), Hit (1), and Stack (2), and train for a total of 3000 epochs. Thus, for example, “BLOCK-012” means that the Push task is trained alone for the first 1000 epochs, followed by a Hit task for the next 1000, and a Stack task for the remaining 1000. This baseline enables us to assess whether the interleaved approach offers improved learning over all possible blocked learning schedules.
V-B Training Settings
All models are implemented using the PyTorch [91] framework, using the same set of hyper-parameters across all baseline methods and our proposed model: a learning rate of $0.001$, a batch size of $100$, a hidden dimension of $4$, and training for $3000$ epochs. Each encoder and decoder module, whether shared or task-specific, consists of two fully connected layers, with ReLU used as the activation function. Optimization is performed using the AdamW optimizer [92], with the AMSGrad variant enabled and default weight decay settings.
Although the SINGLE baseline ends up with a higher number of parameters per task compared to all other models, including ours, we maintain a consistent network architecture across all settings. This design choice ensures that performance differences arise primarily from differences in task interleaving strategies, rather than architectural complexity.
V-C Overall Learning Performance
The initial evaluation of the proposed model compares it to the SINGLE and INTER-RAND baselines. Experiments were conducted using 10 different random seeds, and the average task performance results are presented here. As illustrated in Figure 2, the proposed model INTER-LP reduces prediction error faster than the other two baselines. This indicates that the LP-based task selection with inter-task skill transfer produces an interleaved task schedule that leads to superior overall performance compared random selection and independent learning with no skill transfer. Furthermore, it is important to note that not only does interleaving enhance performance, but simultaneously learning multiple tasks is also beneficial, as tasks can mutually support each other and facilitate positive information transfer. This is evidenced by the INTER-RAND baseline showing improvement compared to the single-task baseline.
@@ V-C Overall Learning Performance
Figure 2: Overall performance of baselines is shown for isolated task learning (SINGLE), proposed architecture but random task selection (INTER-RAND) and proposed architecture with LP-based task selection model (INTER-LP. As can be seen proposed model (INTER-LP) converges faster than the baselines. Even though the random task selection is slightly better than the single task learning, it could not surpass learning with LP-based task selection
Task-wise Analysis. To better understand the learning dynamics, we examine the performance of individual tasks throughout the learning process. As illustrated in Figure 3, the performance on both the Push and Stack tasks improves when all tasks are learned simultaneously, with a marked enhancement observed when using the proposed model. Notably, the Stack task exhibits a significant performance gap between our proposed model, INTER-LP and the other two baselines. This discrepancy likely arises from the Stack task’s greater complexity compared to the other tasks, demonstrating that single-task learning is insufficient for handling such complexity. Hit task, on the other hand, receives a brief early learning improvement during iterations of $300-1000$ and does not show as large gains as the other two tasks. Although the INTER-RAND baseline utilizes a multi-task learning approach, it results in little to no improvement when interleaving is not performed suitably.

### v2

@@ 4.2 Task Arbitration
0: Number of tasks $m$, number of epochs $R$, exploration rate $\epsilon=0.1$
for epoch =1=1= 1 to $R$ do
  Compute score $s^{\mathcal{T}_{i}}$ for tasks $\{\mathcal{T}_{i}\}_{i=1}^{m}$
@@
  Generate a random number $r$ uniformly from [0,1][0,1][ 0 , 1 ]
@@
   $\mathcal{T}_{*}\leftarrow$ Random task from $\{\mathcal{T}_{1},\mathcal{T}_{2},\ldots,\mathcal{T}_{m}\}\setminus\{\mathcal{T}_{i^{\prime}}\}$
@@
   $\mathcal{T}_{*}\leftarrow\mathcal{T}_{i^{\prime}}$
@@
  Perform training on the selected task $\mathcal{T}_{*}$
@@
    Learning Progress (LP) Based Task Selection LP of a task is computed based on the recent history of the error signal pertaining to it: the error in the last $L$ ($L=5$ in the current study) steps linearly regressed against time and negative of its slope is taken as the LP value. To be concrete, let $E^{\mathcal{T}_{i}}_{t}$ be the error at time step $t$ for task $\mathcal{T}_{i}$. Consider the set of errors $\{E^{\mathcal{T}_{i}}_{t-L+1},E^{\mathcal{T}_{i}}_{t-L+2},\ldots,E^{\mathcal{T}_{i}}_{t}\}$. Then, a linear model is fit to these points to estimate the slope $\beta^{\mathcal{T}_{i}}$. If $\beta^{\mathcal{T}_{i}}<0$, the error is decreasing, and the task is making positive learning progress. If $\beta^{\mathcal{T}_{i}}\geq 0$, it suggests that the error has plateaued or is increasing, implying no recent improvement. Under the LP-based interleaving scheme, the proposed task switching mechanism computes the learning progress $LP^{\mathcal{T}_{i}}$ for each task $\{\mathcal{T}_{i}\}_{i=1}^{m}$. Then it selects the task with the highest LP to be trained next. In other words, tasks that are currently showing rapid improvement receive more training time, while tasks that have stalled or regressed receive less immediate attention. Formally, if $\beta^{\mathcal{T}_{i}}$ is the slope for task $\mathcal{T}_{i}$, then $\mathcal{T}_{*}$ is chosen in two steps:
@@
      ${LP^{\mathcal{T}_{i}}}=\begin{cases}\lvert\beta^{\mathcal{T}_{i}}\rvert,&{\text{if }}\ \beta^{\mathcal{T}_{i}}<0\\ {0,}&{\text{otherwise.}}\end{cases}$
@@
    Energy Consumption Modulated Learning Progress (EMLP) Based Task Selection. While learning progress alone can guide the model toward tasks showing improvement, it does not consider the computational or energetic cost associated with training each task. In scenarios where energy efficiency is a concern, we incorporate energy consumption into the interleaving strategy. We define energy consumption $EC^{\mathcal{T}_{i}}$ for task $\mathcal{T}_{i}$ as the total output of all neurons belonging to it $\mathcal{A}^{\mathcal{T}_{i}}$ , cumulated over the last $L$ training steps:
    $EC^{\mathcal{T}_{i}}=\sum_{j=(t-L)}^{t}{\mathcal{A}}^{\mathcal{T}_{i}}_{j}.$
    This could be measured in terms of computational operations, memory usage, or any proxy for energy expenditure. Lower $EC^{\mathcal{T}_{i}}$ values mean the task has recently been trained efficiently, while higher values indicate that the task has been relatively costly in terms of energy. However, using energy consumption only as the interleaving guidance, the network will only choose the tasks that produce low activation energy without taking the learning performance of the tasks into consideration, which can lead to worsened overall multi-task learning performance (see Figure 8(a)). To consider the cost of learning, LP is discounted by a positive constant $k$ that controls the sensitivity of the combined criteria to energy consumption:
    $s^{\mathcal{T}_{i}}=e^{(k\cdot\tilde{LP}^{\mathcal{T}_{i}})}/\tilde{EC}^{\mathcal{T}_{i}}.$ (9)
    Note that, to combine LP and EC properly, they are first scaled between [0,1][0,1][ 0 , 1 ], before the combined score is computed, simply by using a min-max scaler. The $s^{\mathcal{T}_{i}}$ can be interpreted as follows: When $k$ approaches to zero, $e^{(k\cdot\tilde{LP}^{\mathcal{T}_{i}})}\approx 1$, thus the score is primarily influenced by $\tilde{EC}^{\mathcal{T}_{i}}$, even if the task has strong LP. This discourages spending too many ‘energy resources’ on a single, costly task, even if it has high learning gains. On the other hand, when $k$ is high, $e^{(k\cdot\tilde{LP}^{\mathcal{T}_{i}})}$ becomes dominant, thus EC becomes negligible, approximating the previous LP-only selection. Under the EC-modulated LP-based interleaving strategy, we compute combined scores for each task $\mathcal{T}_{i}$ as given above. The next task to train is the one with the highest combined score:
@@
    This ensures a balance between quick gains in task performance and maintaining overall energy efficiency. Tasks with high learning progress but also low recent energy costs are favored, while tasks that have become too energy-intensive or are not showing improvement receive less immediate focus. By integrating these interleaving strategies, we encourage a more dynamic, bio-inspired learning schedule.
5 Experiments and Results
To evaluate the effectiveness of our proposed Interleaved Multitask Learning with Energy Modulated Learning Progress (IMTL-EMLP), a series of experiments are conducted with a simulated robot tasked with learning the effect of its actions in different task settings. In the experiments, we focus on learning efficiency and final task accuracy as performance measures and asses the knowledge transfer among tasks.
The proposed IMTL framework introduces Computational Energy modulated LP (EMLP) for dynamic task selection, which includes LP-based selection as a special case. We begin by evaluating the LP-based approach (IMTL-LP), followed by its energy-aware extension (IMTL-EMLP), which incorporates energy consumption into the task arbitration process.
To isolate the effects of task arbitration strategies, we compare IMTL-LP against several baselines: single-task learning (SINGLE), a multi-task learning model with random task selection (IMTL-RAND), and blocked learning setup where tasks are trained sequentially in fixed permutations (BLOCK). Additionally, we examine IMTL-EMLP, which extends IMTL-LP with energy sensitivity, analyzing how varying the energy coefficient $k$ affects the balance between prediction accuracy and computational cost.
All baselines use a consistent model architecture [^2] The actual layer sizes for both single-task and multi-task learning setups are provided in the Appendix 1. to ensure that performance differences stem solely from task arbitration strategies. In addition to learning efficiency analysis, we assess the robustness to model size by comparing performance across low, medium, and high-resourced networks. Finally, two ablation studies dissect the individual and combined contributions of model components and explore object-wise skill transfer between tasks.
5.1 Simulation Environment
@@
In this study, UR10 robotic arm equipped with a Franka Panda end effector is simulated via the PyBullet engine [103]. We use a tabletop environment to evaluate multi-task learning through effect prediction tasks. A set of basic objects are presented on the table for the robot to interact with during the tasks. Namely, a sphere ($r=3\,\text{cm}$), a cube ($l=6\,\text{cm}$), a square prism ($w=6\,\text{cm},\,h=12\,\text{cm}$), and a cylinder ($r=3\,\text{cm},\,h=18\,\text{cm}$) are used. The latter two objects can be presented in either vertical or horizontal orientation, thus making it six total object configurations for interaction. In each interaction, the object is randomly placed on the table.
5.2 Effect Prediction Tasks
For testing the proposed model, we have defined three tasks for the robotic agent as to predict the effect of its actions in three different settings as described next.
@@ 5.2 Effect Prediction Tasks
Push task. A single object is randomly placed on the table. The robot applies a “pushing” action with its end effector to the object’s center of mass (CoM), with an angle $\theta$ chosen between [0,180][0,180][ 0 , 180 ] degrees. The state of the object is taken as its pose defined by the Cartesian coordinates (x, y, z) and orientation (with Euler angles $\phi_{x},\phi_{y},\phi_{z}$). The orientation angles are represented with sine cosine pairs to avoid discontinuous input for the neural networks that process the object state. Thus, the object is represented as a 9-dimensional vector:
$\mathbf{s}=[x,y,z,\sin(\phi_{x}),\cos(\phi_{x}),\sin(\phi_{y}),\cos(\phi_{y}),\sin(\phi_{z}),\cos(\phi_{z})].$
@@
The effect is defined as the change between the object’s pre-action and post-action states, captured using the simulation environment:
$\mathbf{e}=[\Delta(x),\Delta(y),\Delta(z),\Delta(\sin(\phi_{x})),\Delta(\cos(\phi_{x})),$
$\Delta(\sin(\phi_{y})),\Delta(\cos(\phi_{y})),\Delta(\sin(\phi_{z})),\Delta(\cos(\phi_{z}))].$
Figure 2: Visual examples of tasks executed in the PyBullet simulation environment. Arrows in Push (a) and Hit (b) show the trajectory of the object after the action is executed. For Stack (c), arrows show the end-effector trajectory while executing the stacking action.
@@
Hit Task. The second task is a dynamic variation of the Push task. The setup is the same, but the pushing action is applied with twice the velocity, resulting in less predictable object behavior due to possible multiple bouncing off the walls. As we can see in Figures 2(b), after the vertical cylinder is hit by the robotic arm, it falls over and bounces off from the walls of the table, rolling in the opposite direction of the executed hit action. Here, the tabletop environment and the definitions of $\mathbf{s}$, $\mathbf{a}$, and $\mathbf{e}$ remain identical to the Push task.
@@
Stack Task. In this task, the goal is to place one object on top of another, forming a possible stable stack. This task involves two objects: a ‘picked’ object and a ‘target’ object. As a result, the state includes the positional and orientation information for both objects, doubling the state information compared to the Push and Hit tasks:
$\mathbf{s}=[\mathbf{s}_{p},\mathbf{s}_{t}],$
@@
where $\mathbf{s}_{p}$ and $\mathbf{s}_{t}$ are the 9-dimensional states of the picked and target objects, denoted as $o_{p}$ and $o_{t}$, respectively. Given the set of six object configurations in the study, this task results in 6×6=366$\times$ 6=366 × 6 = 36 possible object pair combinations. The action is represented by concatenating the one-hot encoding of the picked and the target object:
$\mathbf{a}=[\text{onehot}(o_{p}),\text{onehot}(o_{t})].$
Similarly, the effect is computed as the change in the objects pre-action and post-action states:
$\mathbf{e}=[\mathbf{e}_{p},\mathbf{e}_{t}].$
5.3 Compared Models and Baselines
Before evaluating the proposed IMTL-EMLP model, we first examine the properties of IMTL-LP, a special case of IMTL-EMLP. We then shift focus to IMTL-EMLP, comparing its performance against both the IMTL-LP model and various baselines. Additionally, we assess different IMTL-EMLP variants with varying energy sensitivity coefficient $k$, to highlight the trade-off between prediction performance and energy consumption.
The baselines used for assessing advantages of our model with LP-based task selection, i.e., IMTL-LP are as follows:
  • 1.
@@ 5.3 Compared Models and Baselines
    IMTL-RAND: The second baseline uses the shared architecture used in our proposed method, without any modification but removes the LP-based task selection mechanism. Instead, tasks are chosen randomly at each training iteration, with each task having an equal chance of being selected. This baseline serves to highlight the impact of our task selection (interleaving) strategy by comparing it against a scenario where the interleaving is uniformly random. This baseline helps isolate the effectiveness of the scheduling algorithm from the benefits of shared parametrization.
  • 3.
@@
    BLOCK: The third and final baseline is implemented using the same shared network architecture but trains tasks in dedicated, uninterrupted blocks rather than interleaving them. To account for any potential effects caused by the order in which tasks are learned, we evaluate all six possible permutations of the three tasks. For each possible order, each task gets a ‘blocked’ learning period before moving to the next one. If there are $R$ training epochs in total, each task gets $R/m$ number of trainings. This baseline enables us to assess whether the interleaved approach offers improved learning over all possible blocked learning schedules.
For the experiments with the proposed model IMTL-EMLP, we compare it with different variants of it as follows, in addition to the baselines explained above and proposed IMTL-LP model:
  • 1.
    IMTL-EMLP-K: To explore how sensitivity coefficient $k$ affects the learning performance and energy usage, we compare IMTL-EMLP models trained with varying $k$ values: IMTL-EMLP-K=0.4, IMTL-EMLP-K=0.7, IMTL-EMLP-K=1, IMTL-EMLP-K=1.2.
5.4 Network Architecture and Initialization
[…]
@@ 5.4 Network Architecture and Initialization
Figure 3: Overall performance of baselines is shown for isolated task learning (SINGLE), proposed architecture but random task selection (IMTL-RAND) and proposed architecture with LP-based task selection model (IMTL-LP). As can be seen proposed model (IMTL-LP) converges faster than the baselines. Even though the random task selection is slightly better than the single task learning, it could not surpass learning with LP-based task selection.
5.5 LP-based Interleaved Learning (IMTL-LP) Results
The initial evaluation compares the proposed model with no energy modulation, i.e., IMTL-LP to SINGLE and IMTL-RAND baselines. Experiments were conducted using 10 random seeds, and the average task performance results are presented here. As illustrated in Figure 3, the proposed model IMTL-LP reduces prediction error faster than the other two baselines. This indicates that the LP-based task selection with inter-task skill transfer produces an interleaved task schedule that leads to improved overall performance compared to random selection and independent learning with no skill transfer. Furthermore, it is important to note that not only does interleaving enhance performance, learning multiple tasks is also shown to be beneficial, as tasks can mutually support each other and facilitate positive information transfer. This is evidenced by the IMTL-RAND baseline showing improvement compared to the single-task baseline.